OpenStudy (anonymous):

Session75 Recitation Part C

OpenStudy (anonymous):

I keep coming up with \[\int\limits_{}^{}\frac{(2x+2)dx}{(4x^2+1)^2} = \frac{x^2+2x}{1+4x^2} + \frac{arctan(2x)}{2} + C\] but online integration calculator returns \[\frac{arctan(2x)}{2} + \frac{4x-1}{16x^2+4} +C \] So I have the arctan(2x)/2 in common, but my other part doesn't match. I first substitute u = 2x, du = 2dx... \[\frac{(u+2)du}{2(u^2+1)^2}\] And then use trig substitution to set... \[u = \tan(\theta), du = \sec^2(\theta)d\theta\]

OpenStudy (anonymous):

Bah, simple algebra mistakes. CLOSED.